There are 4 red pens, 3 blue pens, 2 black pen, and 4 green pens in a desk drawer. You randomly choose one. Find each probability as a FRACTION in SIMPLEST FORM.
1
Expert's answer
2020-12-17T13:49:28-0500

"P(Pen~is~red) = \frac{4}{4+3+2+4}=\frac{4}{13}\\
P(Pen~is~blue) = \frac{3}{4+3+2+4}=\frac{3}{13}\\
P(Pen~is~black) = \frac{2}{4+3+2+4}=\frac{2}{13}\\
P(Pen\space is\space green) = \frac{4}{4+3+2+4}=\frac{4}{13}"